### v1

# Anderson localization for the multi-frequency quasi-periodic CMV matrices and quantum walks

Abstract
In this paper, we establish Anderson localization for the CMV matrices with multi-frequency analytic quasi-periodic Verblunsky coefficients in the regime of the positive Lyapunov exponent. As an application, we further derive the Anderson localization for the multi-frequency analytic quasi-periodic quantum walks. We extend the results of Wang and Damanik [25] for one-frequency quasi-periodic CMV matrices to multi-frequency case.
keywords:
Quasi-periodic CMV matrix; Multi-frequency; Anderson localization; Quantum walks; Gesztesy-Zinchenko cocycle
MSC:
[2010] 37A30 , 42C05 , 70G60
journal: arXiv
1 Introduction and the statement of main result
The problem of Anderson localization (AL) for CMV matrices and quantum walks has received attention recently. Zhu [29] showed both Anderson localization and strong dynamical localization for random CMV matrices. Wang-Damanik [25] formulated AL for quasiperiodic CMV matrices and quantum walks whose Verblunsky coefficients are defined by shift. Cedzich-Werner [6] obtained AL for a class of electric quantum walks and skew-shift CMV matrices. Lin-Piao-Guo [21] proved AL for CMV matrices with Verblunsky coefficients defined by skew-shift. They extended the localization results for Schrödinger operators in [2] and [4] to CMV matrices respectively. In the sequel, Lin-Guo-Piao [20] proved AL for CMV matrices whose Verblunsky coefficients are given by a hyperbolic toral automorphism. They extended the localization results for Schrödinger operators in [2] and [4] to CMV matrices.
To the best of our knowledge, the localization problem of the multi-frequency quasi-periodic CMV matrices still remains unknown. This is precisely the issue that the present paper focuses on. We are strongly motivated by the papers [7, 2, 3, 17, 28, 15, 16] whose are about AL and related spectral analysis for the multi-frequency quasi-periodic Schrödinger operators. Generally speaking, in the spectral theory of quasi-periodic Schrödinger operators, we observe that the analysis of AL for multi-frequency operators is more complicated than the one-frequency case. The most obvious obstruction is the elimination of double resonances along the orbit; see the references [1, 7]. Our study indicates that the CMV matrix situation is also similar. Indeed, there is a close relation between the spectral theories of CMV operators and one-dimensional Schrödinger operators. The detailed review for the relationship one can refer to the references [21, 25, 22, 23]. But it is necessary to emphasize that working out the CMV analog of a Schrödinger result is not straightforward.
For the one-frequency quasi-periodic CMV matrices, Wang-Damanik [25] got the elimination of double resonances along the orbit directly by using the frequency estimate [2, Lemma 6.1]. However, to our knowledge, there is no such estimate for multi-frequency CMV matrices. For this issue, we apply the semialgebraic sets arguments [1, Corollary 9.7 and Lemma 9.9]. Therefore, one of the key steps is to construct a proper semialgebraic set and do the corresponding decomposition. The other key steps include mainly the avalanche principle (AP) and the large deviation theorem (LDT).
For recent work on other spectral properties of CMV matrices, such as the regularity of Lyapunov exponent, absolutely continuous spectrum, etc., there are also fruitful results and one can consult the interesting papers [10, 11, 12, 19, 26, 27].
The main result of this paper can be stated as follows:
Theorem 1.1.
The multi-frequency quasi-periodic CMV matrices $\mathcal{E}$ whose Verblunsky coefficients are defined by a shift display Anderson localization in the regime of positive Lyapunov exponents for a fixed phase vector $x_{0}$ and almost all frequency vectors satisfying the standard Diophantine condition; that is, $\mathcal{E}(x_{0})$ has pure point spectrum with exponentially decaying eigenfunctions, where $\mathcal{E}(x_{0})$ denotes the extended CMV matrix with phase $x_{0}$.
The rest of the present paper is organized as follows: we provide a quick introduction to the orthogonal polynomials on the unit circle (OPUC) and the settings in Section 2. Section 3 contains some basic tools which are helpful to the following study. Section 4 addresses the proof of the main theorem. Finally, we show AL for the analytic multi-frequency quasi-periodic quantum walks in Section 5.
2 Preliminaries
Let us recall some preliminaries of OPUC that can be seen for more details in [22]. Let $\mathbb{D}=\{z:|z|<1\}$ be the open unit disk in $\mathbb{C}$ and $\mu$ be a nontrivial probability measure on $\partial\mathbb{D}=\{z:|z|=1\}$. Therefore, the functions $1,z,z^{2},\ldots$ are linearly independent in the Hilbert space $L^{2}(\partial\mathbb{D},d\mu)$. Let $\Phi_{n}(z)$ be the monic orthogonal polynomials; that is, $\Phi_{n}(z)=P_{n}[z^{n}]$, $P_{n}\equiv$ projection onto $\{1,z,z^{2},\ldots,z^{n-1}\}^{\bot}$ in $L^{2}(\partial\mathbb{D},d\mu)$. Then the orthonormal polynomials are
$\varphi_{n}(z)=\frac{\Phi_{n}(z)}{\|\Phi_{n}(z)\|_{\mu}},$
where $\|\cdot\|_{\mu}$ denotes the norm of $L^{2}(\partial\mathbb{D},d\mu)$.
For any polynomial $Q_{n}(z)$ of degree $n$, one can define its Szegő dual $Q_{n}^{*}(z)$ by
$Q_{n}^{*}(z)=z^{n}\overline{Q_{n}(1/\overline{z})}.$
Then the Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}$ ($\alpha_{n}\in\mathbb{D}$) obey the equation
$\Phi_{n+1}(z)=z\Phi_{n}(z)-\overline{\alpha}_{n}\Phi_{n}^{*}(z),$ (2.1)
which is called Szegő recursion.
Orthonormalizing $1,z,z^{-1},z^{2},z^{-2},\cdots$, one can get a CMV basis $\{\chi_{j}\}_{j=0}^{\infty}$ of $L^{2}(\partial\mathbb{D},d\mu)$, and the matrix representation of multiplication by $z$ relative to the CMV basis gives rise to the CMV matrix $\mathcal{C}$,
$\mathcal{C}_{ij}=\langle\chi_{i},z\chi_{j}\rangle.$
Then $\mathcal{C}$ has the form
$\mathcal{C}=\left[\begin{array}[]{ccccccc}\overline{\alpha}_{0}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ \rho_{0}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&\ddots&\ddots&\ddots\end{array}\right],$
where $\rho_{j}=\sqrt{1-|\alpha_{j}|^{2}}$, $j\geq 0$. Similarly, we can get the extended CMV matrix
$\mathcal{E}=\left[\begin{array}[]{cccccccc}\ddots&\ddots&\ddots&&&&&\\ &-\overline{\alpha}_{0}\alpha_{-1}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ &-\rho_{0}\alpha_{-1}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &&\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&&\ddots&\ddots&\ddots\end{array}\right],$
which is a special five-diagonal doubly-infinite unitary matrix in the standard basis of $L^{2}(\partial\mathbb{D},d\mu)$ according to [22, Subsection 4.5] and [23, Subsection 10.5].
In this paper, we consider a sequence of Verblunsky coefficients which are generated by an analytic function $\alpha(\cdot):\mathbb{T}^{d}\rightarrow\mathbb{D}$, $\alpha_{n}(x)=\alpha(T^{n}x)=\alpha(x+n\omega)$, where $T$ is an invertible map of the form $Tx=x+\omega$, $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$, $x,\omega\in\mathbb{T}^{d}$ are phase and frequency, respectively. Moreover, the frequency vector $\omega\in\mathbb{T}^{d}$ satisfies the standard Diophantine condition
$\|k\cdot\omega\|\geq\frac{p}{|k|^{q}}$ (2.2)
[…]
Assume that the sampling function $\alpha(x)$ satisfies that
[…]
and it can extend complex analytically to the region
$\mathbb{T}^{d}_{h}:=\{x+iy:x\in\mathbb{T}^{d},y\in\mathbb{R}^{d},|y|<h\}$
for some constant $h>0$.
Then the Szegő recursion is equivalent to
$\rho_{n}(x)\varphi_{n+1}(z)=z\varphi_{n}(z)-\overline{\alpha}_{n}(x)\varphi_{n}^{*}(z),$ (2.4)
where $\rho_{n}(x)=\rho(x+n\omega)$ and $\rho(x)=(1-|\alpha(x)|^{2})^{1/2}$. Applying Szegő dual to both sides of (2.4), one can obtain that
$\rho_{n}(x)\varphi_{n+1}^{*}(z)=\varphi_{n}^{*}(z)-\alpha_{n}(x)z\varphi_{n}(z).$ (2.5)
Then one can get the matrix form of (2.4) and (2.5), i.e.,
$\left[\begin{array}[]{c}\varphi_{n+1}\\ \varphi_{n+1}^{*}\end{array}\right]=S(\omega,z;x+n\omega)\left[\begin{array}[]{c}\varphi_{n}\\ \varphi_{n}^{*}\end{array}\right],$
where
$S(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}z&-\overline{\alpha}(x)\\ -\alpha(x)z&1\end{array}\right].$
Since $\det S(\omega,z;x)=z$, one always prefers to study the matrix
$M(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\alpha}(x)}{\sqrt{z}}\\ -\alpha(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right]\in\mathbb{S}\mathbb{U}(1,1).$
This is called the Szegő cocycle map. It is easy to verify that $\mathrm{det}M(\omega,z;x)=1$. Then the monodromy matrix (or $n$-step transfer matrix) is defined by
$M_{n}(\omega,z;x)=\prod_{j=n-1}^{0}M(\omega,z;x+j\omega)$ (2.6)
According to (2.6), it is obvious that
$M_{n_{1}+n_{2}}(\omega,z;x)=M_{n_{2}}(\omega,z;x+n_{1}\omega)M_{n_{1}}(\omega,z;x)$
and
$\log\|M_{n_{1}+n_{2}}(\omega,z;x)\|\leq\log\|M_{n_{1}}(\omega,z;x)\|+\log\|M_{n_{2}}(\omega,z;x+n_{1}\omega)\|.$ (2.7)
Let
$u_{n}(\omega,z;x):=\frac{1}{n}\log\|M_{n}(\omega,z;x)\|$ (2.8)
and
[…]
Integrating the inequality (2.7) with respect to $x$ over $\mathbb{T}^{d}$, we have that
$L_{n_{1}+n_{2}}(\omega,z)\leq\frac{n_{1}}{n_{1}+n_{2}}L_{n_{1}}(\omega,z)+\frac{n_{2}}{n_{1}+n_{2}}L_{n_{2}}(\omega,z).$
This implies that
$L_{n}(\omega,z)\leq L_{m}(\omega,z)\quad\mathrm{if}\quad m<n,\;m|n$
and
$L_{n}(\omega,z)\leq L_{m}(\omega,z)+C\frac{m}{n}\quad\mathrm{if}\quad m<n.$
As we know, the transformation $x\rightarrow x+\omega$ is ergodic for any irrational $\omega\in\mathbb{T}^{d}$. Notice that (2.7) implies that $\log\|M_{n}(\omega,z;x)\|$ is subadditive. Thus, according to Kingman’s subadditive ergodic theorem, the limit
$L(\omega,z)=\lim_{n\rightarrow\infty}L_{n}(\omega,z)$ (2.10)
exists. This is called the Lyapunov exponent. Throughout this paper, we let $\gamma$ be the lower bound of the Lyapunov exponent. On the other hand, Furstenberg-Kesten theorem indicates that the limit also exists for $a.e.\;x$:
$\lim_{n\rightarrow\infty}u_{n}(\omega,z;x)=\lim_{n\rightarrow\infty}L_{n}(\omega,z)=L(\omega,z).$ (2.11)
3 Basic tools
In this section, we require some useful lemmas which could have applications in the following research. To begin with, we introduce some useful notations.
Definition 1.
[5, Definition 2.3] For any positive numbers $a,b$ the notation $a\lesssim b$ means $Ca\leq b$ for some constant $C>0$. By $a\ll b$ we mean that the constant $C$ is very large. If both $a\lesssim b$ and $a\gtrsim b$, then we write $a\asymp b$.
3.1 Avalanche principle
Lemma 3.1.
[14, Proposition 2.2] Let $A_{1},\ldots,A_{m}$ be a sequence of arbitrary unimodular $2\times 2$-matrices. Suppose that
$\min_{1\leq j\leq m}\|A_{j}\|\geq\mu\geq m$ (3.1)
and
$\max_{1\leq j<m}\big{[}\log\|A_{j+1}\|+\log\|A_{j}\|-\log\|A_{j+1}A_{j}\|\big{]}<\frac{1}{2}\log\mu.$ (3.2)
Then
$\Big{|}\log\|A_{m}\cdots A_{1}\|+\sum_{j=2}^{m-1}\log\|A_{j}\|-\sum_{j=1}^{m-1}\log\|A_{j+1}A_{j}\|\Big{|}<C_{A}\frac{m}{\mu},$ (3.3)
where $C_{A}$ is an absolute constant.
3.2 Large deviation theorem
Now we turn attention to proving the large deviation theorem which is the most basic tool in the theory of localization for CMV matrices. To obtain LDT, we need the following lemma.
Lemma 3.2.
[14, Proposition 9.1] Let $d$ be a positive integer. Suppose $u:D(0,2)^{d}\rightarrow[-1,1]$ is subharmonic in each variable; i.e., $z_{1}\rightarrow u(z_{1},z_{2},\ldots,z_{d})$ is subharmonic for any choice of $(z_{2},\ldots,z_{d})\in D(0,2)^{d-1}$ and similarly for each of the other variables. Assume furthermore that for some $n\geq 1$
$\sup_{\theta\in\mathbb{T}^{d}}|u(\theta+\omega)-u(\theta)|<\frac{1}{n}.$ (3.4)
Then there exist $\sigma>0$, $\tau>0$, and $c_{0}$ only depending on $d$ and $\varepsilon_{1}$ such that
$\mathrm{mes}\{\theta\in\mathbb{T}^{d}:|u(\theta)-\langle u\rangle|>n^{-\tau}\}<\exp(-c_{0}n^{\sigma}).$ (3.5)
[…]
Based on the above lemma, we get the large deviation estimate.
Theorem 3.3.
Suppose $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. There exist $\sigma=\sigma(p,q)$, $\tau=\tau(p,q)$, $\sigma,\tau\in(0,1)$, $C_{0}=C_{0}(p,q,h)$ such that for $n\geq 1$ one has that
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x)\|-nL_{n}(\omega,z)|>n^{1-\tau}\}<\exp(-C_{0}n^{\sigma}).$ (3.6)
Proof.
Fix some dimension $d$ and $z$. For any $x\in D(0,2)^{d}$, define
$u_{n}(x)=\frac{1}{n}\log\|M_{n}(\omega,z;x)\|.$
Then $u_{n}$ is a continuous subharmonic function which is bounded by $1$ in $D(0,1)^{d}$. In addition, $u_{n}$ satisfies the conditions in the above lemma. Thus, (3.6) is an immediate consequence of (3.5). ∎
3.3 Estimate for the Lyapunov exponent
Lemma 3.4.
Assume $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. Then for any $n\geq 2$,
$0\leq L_{n}(\omega,z)-L(\omega,z)<C\frac{(\log n)^{1/\sigma}}{n},$
where $C=C(p,q,z,\gamma)$ and $\sigma$ is as in LDT.
Proof.
Obviously, $0\leq L_{n}(\omega,z)\leq C(z)$ for all $n$. Let $k$ be a positive integer such that $k\gamma>16C(z)$. Given $n>10$, let $l=[C_{1}(\log n)^{1/\sigma}]$ with some large $C_{1}$. Consider the integers $l_{0},2l_{0},\ldots,2^{k}l_{0}$. Then there exists some $0\leq j<k$ such that (with $l_{j}=2^{j}l_{0}$)
$L_{l_{j}}(z)-L_{l_{j+1}}(z)<\frac{\gamma}{16}.$ (3.7)
If not, then $C(z)>L_{l_{0}}(z)-L_{2^{k}l_{0}}(z)\geq\frac{k\gamma}{16}>C(z)$, which is a contradiction. Let $l=2^{j}l_{0}$ with the choice of $j$ satisfying (3.7).
It is not difficult to check that $M(\omega,z;x)$ is conjugate to an $SL(2,\mathbb{R})$ matrix $T(\omega,z;x)$ through the matrix
$Q=-\frac{1}{1+i}\left[\begin{array}[]{cc}1&-i\\ 1&i\end{array}\right]\in\mathbb{U}(2);$
that is,
$T(\omega,z;x)=Q^{*}M(\omega,z;x)Q\in SL(2,\mathbb{R}),$
where “*” stands for the conjugate transpose of a matrix.
Now we apply the avalanche principle (AP) to the matrices $B_{j}=T_{l}(\omega,z;x+(j-1)l\omega)$ for $1\leq j\leq m=[n/l]$ with $\mu=\exp(l\gamma/2)$. Note that $\mu>n^{2}$ if $C_{1}>4/\gamma$. According to (LDT), the phases $x$ such that
$\min_{1\leq j\leq m}\|B_{j}\|\geq\exp(lL_{l}(\omega,z)-l^{1-\tau})>\exp(l\gamma/2)=\mu>n^{2}$
form a set $\mathcal{G}_{1}$ whose measure is not exceeding
$m\exp(-C_{0}l^{\sigma})\leq n\exp(-C_{0}(2^{j}l_{0})^{\sigma})<n^{-2}$
provided $C_{1}>n^{1/\sigma}$.
Furthermore, combining (3.7) and (LDT),
$\max_{1\leq j<m}\Big{|}\log\|B_{j+1}\|+\log\|B_{j}\|-\log\|B_{j+1}B_{j}\|\Big{|}$ $\leq 2l(L_{l}(z)+\frac{\gamma}{32})-2l(L_{2l}(z)-\frac{\gamma}{32})$
$=2l[(L_{l}(z)-L_{2l}(z))+\frac{\gamma}{16})]$
$\leq\exp(l\gamma/4)=\frac{1}{2}\log\mu$
up to a set $\mathcal{G}_{2}$ of $x$ of measure not exceeding
$2m\exp(-C_{0}l^{\sigma})\leq 2n\exp(-C_{0}(l)^{\sigma})<n^{-2}$
if $C_{1}>n^{1/\sigma}$. Let $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$. Then $\mathrm{mes}\mathcal{G}<2n^{-2}$ and (3.1), (3.2) hold for all $x\in\mathcal{G}$.
Based on (AP), we have that
$\Big{|}\log\|B_{m}\cdots B_{1}\|+\sum_{j=2}^{m-1}\log\|B_{j}\|-\sum_{j=1}^{m-1}\log\|B_{j+1}B_{j}\|\Big{|}<C\frac{n}{\mu}<Cn^{-1}$
for all $x\in\mathcal{G}$. Equivalently, one can obtain that
$\Big{|}\log\|T_{lm}(\omega,z;x)+\sum_{j=2}^{m-1}\log\|T_{l}(\omega,z;x+jl\omega)\|-\sum_{j=1}^{m-1}\log\|T_{l}(\omega,z;x+(j+1)l\omega)T_{l}(\omega,z;x+jl\omega)\|\Big{|}<Cn^{-1}$
for all $x\in\mathcal{G}$.
Similarly, for $\log\|T_{lm}(\omega,z;x+lm\omega)\|$ and $\log\|T_{lm}(\omega,z;x)\|$, the phases $x$ such that
$\Big{|}\|\log\|T_{2lm}(\omega,z;x)\|-\|\log\|T_{lm}(\omega,z;x+lm\omega)\|-\|\log\|T_{lm}(\omega,z;x)\|$
$+\|\log\|T_{l}(\omega,z;x+lm\omega)\|+\|\log\|T_{l}(\omega,z;x+(m-1)l\omega)\|$
$-\|\log\|T_{l}(\omega,z;x+lm\omega)T_{l}(\omega,z;x+(m-1)l\omega)\|\Big{|}\leq\frac{C}{n}$ (3.8)
form a set with measure not exceeding $Cn^{-2}$.
Since
$\big{|}\|\log\|T_{n}(\omega,z;x)\|-\log\|T_{lm}(\omega,z;x)\|\big{|}\leq C(\alpha_{n},\alpha_{lm},z)n$ (3.9)
and
$\big{|}\|\log\|T_{l}(\omega,z;x)\|\big{|}\leq C(z)l,$ (3.10)
we can conclude from (3.8) that
$\big{|}\|\log\|T_{2n}(\omega,z;x)\|-\log\|T_{n}(\omega,z;x+n\omega)\|-\log\|T_{n}(\omega,z;x)\|\big{|}\leq C(\log n)^{1/\sigma}$
up to a set of $x$ not exceeding $Cn^{-2}$ in measure.
Integrating the above inequality over $x$, we have that
$|L_{2n}(z)-L_{n}(z)|\leq C\frac{(\log n)^{1/\sigma}}{n},$
where $C=C(\gamma,z,\omega)$.
We can finally prove this lemma by summing over $2^{k}n$ ∎
3.4 Green’s function and Poisson’s formula
Define the unitary matrices
$\Theta_{n}=\left[\begin{array}[]{cc}\overline{\alpha}_{n}&\rho_{n}\\ \rho_{n}&-\alpha_{n}\end{array}\right].$
Then one can factorize the matrix $\mathcal{C}$ as
$\mathcal{C}=\mathcal{L}_{+}\mathcal{M}_{+},$
where
$\mathcal{L}_{+}=\left[\begin{matrix}\Theta_{0}&~{}&~{}\\ ~{}&\Theta_{2}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right],\quad\mathcal{M}_{+}=\left[\begin{matrix}\mathbf{1}&~{}&~{}\\ ~{}&\Theta_{1}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right].$
Similarly, the extended CMV matrix can be written as
$\mathcal{E}=\mathcal{L}\mathcal{M},$
where
$\mathcal{L}=\bigoplus_{j\in\mathbb{Z}}\Theta_{2j},\quad\mathcal{M}=\bigoplus_{j\in\mathbb{Z}}\Theta_{2j+1}.$
We let $\mathcal{E}_{[a,b]}$ denote the restriction of an extended CMV matrix to a finite interval $[a,b]\subset\mathbb{Z}$, defined by
$\mathcal{E}_{[a,b]}=P_{[a,b]}\mathcal{E}(P_{[a,b]})^{*},$
where $P_{[a,b]}$ is the projection $\ell^{2}(\mathbb{Z})\rightarrow\ell^{2}([a,b])$. $\mathcal{L}_{[a,b]}$ and $\mathcal{M}_{[a,b]}$ are defined similarly.
However, the matrix $\mathcal{E}_{[a,b]}$ will no longer be unitary due to the fact that $|\alpha_{a-1}|<1$ and $|\alpha_{b}|<1$. To solve this issue, we need to modify the boundary conditions briefly. With $\beta,\eta\in\partial\mathbb{D}$, define the sequence of Verblunsky coefficients
$\tilde{\alpha}_{n}=\begin{cases}\beta,\quad&n=a-1;\\ \eta,&n=b;\\ \alpha_{n},&n\notin\{a-1,b\}.\end{cases}$
Denote the extended CMV matrix with Verblunsky coefficients $\tilde{\alpha}_{n}$ by $\tilde{\mathcal{E}}$. Define
$\mathcal{E}_{[a,b]}^{\beta,\eta}=P_{[a,b]}\tilde{\mathcal{E}}(P_{[a,b]})^{*}.$
$\mathcal{L}_{[a,b]}^{\beta,\eta}$ and $\mathcal{M}_{[a,b]}^{\beta,\eta}$ are defined correspondingly. Then $\mathcal{E}_{[a,b]}^{\beta,\eta}$, $\mathcal{L}_{[a,b]}^{\beta,\eta}$ and $\mathcal{M}_{[a,b]}^{\beta,\eta}$ are all unitary.
For $z\in\mathbb{C}$, $\beta,\eta\in\partial\mathbb{D}$, we can define the characteristic determinant of matrix [$\mathcal{E}^{\beta,\eta}_{[a,b]}$,
$\varphi^{\beta,\eta}_{[a,b]}(z):=\det(z-\mathcal{E}^{\beta,\eta}_{[a,b]}),\quad\phi^{\beta,\eta}_{[a,b]}(z):=(\rho_{a}\cdots\rho_{b})^{-1}\varphi^{\beta,\eta}_{[a,b]}(z).$
Note that when $a>b$, $\phi^{\beta,\eta}_{[a,b]}(z)=1$.
Since the equation ~$\tilde{\mathcal{E}}u=zu$ is equivalent to $(z\mathcal{L}^{*}-\mathcal{M})u=0$, the associated finite-volume Green’s functions are as follows:
$G^{\beta,\eta}_{[a,b]}(z)=\big{(}z(\mathcal{L}^{\beta,\eta}_{[a,b]})^{*}-\mathcal{M}^{\beta,\eta}_{[a,b]}\big{)}^{-1},$
$G^{\beta,\eta}_{[a,b]}(j,k;z)=\langle\delta_{j},G^{\beta,\eta}_{[a,b]}(z)\delta_{k}\rangle,\quad j,k\in[a,b].$
According to [18, Proposition 3.8] and [29, Section B.1], for $\beta,\eta\in\partial\mathbb{D}$, the Green’s function has the expression:
$|G^{\beta,\eta}_{[a,b]}(j,k;z)|=\rho_{j}\cdots\rho_{k-1}\Big{|}\frac{\varphi^{\beta,\cdot}_{[a,j-1]}(z)\varphi^{\cdot,\eta}_{[k+1,b]}(z)}{\varphi^{\beta,\eta}_{[a,b]}(z)}\Big{|},\quad a\leq j\leq k\leq b,$
where $\cdot$” stands for the unchanged Verblunsky coefficient.
From [18, Lemma 3.9], if $u$ satisfies ~$\tilde{\mathcal{E}}u=zu$, Poisson’s formula reads
$u(m)=$ $G^{\beta,\eta}_{[a,b]}(a,m;z)\begin{cases}(z\overline{\beta}-\alpha_{a})u(a)-\rho_{a}u(a+1),\quad&a\text{ even}\\ (z\alpha_{a}-\beta)u(a)+z\rho_{a}u(a+1),&a\text{ odd}\end{cases}$
$+G^{\beta,\eta}_{[a,b]}(m,b;z)\begin{cases}(z\overline{\eta}-\alpha_{b})u(b)-\rho_{b}u(b-1),\quad&b\text{ even}\\ (z\alpha_{b}-\eta)u(b)+z\rho_{b-1}u(b-1),&b\text{ odd}\end{cases}$
for $a<m<b$.
Without loss of generality, we restrict the extended CMV matrix to the interval $[0,n-1]\subset\mathbb{Z}$. Let $\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$ be the characteristic determinant of matrix [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}$. It follows from [24, Theorem 2] that the relation between the $n$-step transfer matrix and the characteristic determinant is
$M_{n}(\omega,z;x)=(\sqrt{z})^{-n}\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}\left(\begin{matrix}z\varphi^{\beta,\eta}_{[1,n-1]}&\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\\ z\Big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\Big{)}^{*}&\big{(}\varphi^{\beta,\eta}_{[1,n-1]}\big{)}^{*}\end{matrix}\right).$ (3.11)
Based upon the above analysis, we can get the Green’s function estimate which will be of great significance in the proof of the Anderson localization.
Lemma 3.5.
Assume $L(z)>\gamma$. Then for $n\gg N_{0}(\gamma)$, there exists a set $\Omega\subset\mathbb{T}^{d}$ satisfying
$\mathrm{mes}\Omega<\exp(-C_{0}n^{\sigma}),$ (3.12)
where $C_{0}$ and $\sigma$ are as in (LDT). Furthermore, for any $x$ outside $\Omega$,
$|G_{\Lambda}^{\beta,\eta}(j,k;z)|<e^{-L(z)|j-k|+n^{1-}}$ (3.13)
holds, where $\Lambda$ is one of
$\{[0,n-1],[1,n-1]\}.$
Proof.
Without loss of generality, we assume that $10\leq j\leq k\leq n-$ 1 and $n\gg N_{0}(\gamma)$. Then one can obtain that
$|G^{\beta,\eta}_{[0,n-1]}(j,k;z)|$ $=\rho_{j}\cdots\rho_{k-1}\Big{|}\frac{\varphi^{\beta,\cdot}_{[0,j-1]}(z)\varphi^{\cdot,\eta}_{[k+1,n-1]}(z)}{\varphi^{\beta,\eta}_{[0,n-1]}(z)}\Big{|}$
$\leq\Big{|}\frac{\varphi^{\beta,\cdot}_{[0,j-1]}(z)\varphi^{\cdot,\eta}_{[k+1,n-1]}(z)}{\varphi^{\beta,\eta}_{[0,n-1]}(z)}\Big{|}$
$\lesssim\frac{\|M_{j}(\omega,z;x)\|\|M_{n-k}(\omega,z;T^{k}x)\|}{\|M_{n}(\omega,z;x)\|}$
provided that $\varphi^{\beta,\eta}_{[0,n-1]}(z)\gtrsim\|M_{n}(\omega,z;x)\|$.
Let $\Omega$ be the set of LDT. That is to say,
$\Omega=\{x\in\mathbb{T}^{d}:\big{|}\log\|M_{n}(\omega,z;x)\|-nL_{n}(\omega,z)\big{|}>n^{1-\tau}\}.$
According to the relation between the characteristic determinant and the $n$-step transfer matrix, for $x\notin\Omega$, we get that
$e^{nL_{n}(\omega,z)-n^{1-\tau}}<\|M_{n}(\omega,z;x)\|<e^{nL_{n}(\omega,z)+n^{1-\tau}}.$
From Lemma 3.4, we have that
$L_{n}(\omega,z)<L(\omega,z)+C\frac{(\log n)^{1/\sigma}}{n}.$
Thus,
$e^{nL(\omega,z)+C(\log n)^{1/\sigma}-n^{1-\tau}}<\|M_{n}(\omega,z;x)\|<e^{nL(\omega,z)+C(\log n)^{1/\sigma}+n^{1-\tau}}.$
As a consequence, we obtain that
$|G^{\beta,\eta}_{[0,n-1]}(j,k;z)|$ $\lesssim\frac{\|M_{j}(\omega,z;x)\|\|M_{n-k}(\omega,z;T^{k}x)\|}{\|M_{n}(\omega,z;x)\|}$
$\leq\frac{e^{(n-(k-j))L(z)+2C(\log n)^{1/\sigma}+n^{1-\tau}}}{e^{nL(z)+C(\log n)^{1/\sigma}-n^{1-\tau}}}$
$\leq e^{-L(z)|j-k|+n^{1-}}.$
3.5 Semialgebraic sets
In this section, we recall some basic preliminaries about the semialgebraic sets developed by Bourgain, which are extremely powerful tools to prove the main theorem in the present paper.
Definition 2.
[1, Definition 9.1] A set $\mathcal{S}\subset\mathbb{R}^{n}$ is called semialgebraic if it is a finite union of sets defined by a finite number of polynomial equalities and inequalities. More precisely, let $\mathcal{P}=\{P_{1},\ldots,P_{s}\}\subset R[X_{1},\ldots,X_{n}]$ be a family of real polynomials whose degrees are bounded by $d$. A (closed) semialgebraic set $\mathcal{S}$ is given by an expression
$\mathcal{S}=\bigcup\limits_{j}\bigcap\limits_{l\in\mathcal{L}_{j}}\{R^{n}|P_{l}s_{jl}0\},$ (3.14)
where $\mathcal{L}_{j}\subset\{1,\ldots,s\}$ and $s_{jl}\in\{\geq,\leq,=\}$ are arbitrary. We say that $\mathcal{S}$ has degree at most $sd$, and its degree is the infimum of $sd$ over all representations as in (3.14).
Lemma 3.6.
[1, Proposition 9.2] Let $\mathcal{S}\subset\mathbb{R}^{n}$ be semialgebraic defined in terms of $s$ polynomials of degree at most $d$ as in (3.14). Then there exists a semialgebraic description of its projection onto $\mathbb{R}^{n-1}$ by a formula involving at most $s^{2n}d^{O(n)}$ polynomials of degree at most $d^{O(n)}$. In particular, if $\mathcal{S}$ has degree $B$, then any projection of $\mathcal{S}$ has degree at most $B^{C}$, $C=C(n)$.
Lemma 3.7.
[1, Corollary 9.7] Let $\mathcal{S}\subset[0,1]^{n}$ be semialgebraic of degree $B$ and $\mathrm{mes}_{n}\mathcal{S}<\eta$. Let $\omega\in\mathbb{T}^{n}$ satisfy Diophantine condition and $N$ be a large integer,
$\log B\ll\log N<\log\frac{1}{\eta}.$
Then for any $x_{0}\in\mathbb{T}^{n}$,
[…]
for some $\delta=\delta(\omega)$.
Lemma 3.8.
[1, Lemma 9.9] Let $\mathcal{S}\subset[0,1]^{2n}$ be semialgebraic of degree $B$ and $\mathrm{mes}_{2n}\mathcal{S}<\eta$, $\log B\ll\log\frac{1}{\eta}$. We denote $(\omega,x)\in[0,1]^{n}\times[0,1]^{n}$ the product variable. Fix $\varepsilon>\eta^{\frac{1}{2n}}$. Then there is a decomposition
$\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}$
with $\mathcal{S}_{1}$ satisfying
$\mathrm{mes}(\mathrm{Proj}_{\omega}\mathcal{S}_{1})<B^{C}\varepsilon$ (3.16)
and $\mathcal{S}_{2}$ satisfying the transversality property
$\mathrm{mes}(\mathcal{S}_{2}\cap L)<B^{C}\varepsilon^{-1}\eta^{1/2n}$ (3.17)
for any $n$-dimensional hyperplane $L$ such that $\mathop{\max}\limits_{0\leq j\leq n-1}|\mathrm{Proj}_{L}(e_{j})|<\frac{1}{100}\varepsilon$ (we denote $(e_{0},\ldots,e_{n-1})$ the $\omega$-coordinate vectors).
4 Proof of the main result
For easier reading, we recall the paving property [25] which we will use in the proof of the main result.
For $\Lambda\subset\mathbb{Z}$, we let
$\mathcal{E}_{\Lambda}=R_{\Lambda}\mathcal{E}R_{\Lambda}^{*},$
where $R_{\Lambda}$ is the restriction operator and
$G_{\Lambda}=(z-\mathcal{E}_{\Lambda})^{-1}.$
Recall the resolvent identity,
$G_{\Lambda}=(G_{\Lambda_{1}}+G_{\Lambda_{2}})-(G_{\Lambda_{1}}+G_{\Lambda_{2}})(\mathcal{E}_{\Lambda}-\mathcal{E}_{\Lambda_{1}}-\mathcal{E}_{\Lambda_{2}})G_{\Lambda},$
where $\Lambda$ is a disjoint union $\Lambda=\Lambda_{1}\cup\Lambda_{2}$ provided that the inverses make sense. One of the consequences of the resolvent identity is the following paving property.
We let $I\subset\mathbb{Z}$ be an interval of size $N>n$ such that for each $x\in I$, there is an interval $I^{\prime}\subset I$ of size $n$ satisfying
$\big{\{}y\in I:|x-y|<\frac{n}{10}\big{\}}\subset I^{\prime}$
and
$|G_{I^{\prime}}(n_{1},n_{2})|<e^{-c|n_{1}-n_{2}|}$
for some constant $c>0$ and $n$ sufficiently large. Then we also have that
$|G_{I}(n_{1},n_{2})|<e^{-\frac{c}{2}|n_{1}-n_{2}|}.$
For the proof of this statement, one can consult part (IV) of [2].
Proof of Theorem 1.1 Fix $n_{0}$ and consider the property
$|G_{[0,n_{0}-1]}^{\beta,\eta}(m_{1},m_{2};z)|<e^{-\gamma|m_{1}-m_{2}|+n_{0}^{1-}}$ (4.1)
for all $10\leq m_{1},m_{2}\leq n_{0}-$ 1. Replace (4.1) by the condition
$\mathop{\sum}\limits_{0\leq m_{1},m_{2}\leq n_{0}-1}$ $e^{2\gamma|m_{1}-m_{2}|}\big{|}\mathrm{det}[(m_{1},m_{2})-\mathrm{minor}\,\,\mathrm{of}\,\,(z-\mathcal{E}_{[0,n_{0}-1]}^{\beta,\eta}(x,\omega))]\big{|}^{2}$
$<e^{2n_{0}^{1-}}\big{|}\mathrm{det}(z-\mathcal{E}_{[0,n_{0}-1]}^{\beta,\eta}(x,\omega))\big{|}^{2}.$ (4.2)
For the condition (4.1), its left-hand side is a polynomial in ($\mathrm{Re}\alpha(x),\mathrm{Im}\alpha(x),\mathrm{Re}z,\mathrm{Im}z$). Since $\mathrm{Re}\alpha(x)$ and $\mathrm{Im}\alpha(x)$ are real analytic functions, then we can replace them by the trigonometric polynomials
$f(x)=\mathop{\sum}\limits_{|k|\leq Cn_{0}}f_{k}e^{i\langle k,x\rangle}\quad\mathrm{and}\quad g(x)=\mathop{\sum}\limits_{|k|\leq Cn_{0}}g_{k}e^{i\langle k,x\rangle}$
respectively with error less than $e^{-\tilde{C}n_{0}^{2}}$ ($\tilde{C}$ a constant), where $|f_{k}|,|g_{k}|<\exp(-r|k|)$, $r$ is a constant, $C$ is a sufficiently large constant. Then we can obtain a polynomial inequality
$P(\cos\omega_{1},\cdots,\cos\omega_{d},\sin\omega_{1},\cdots,\sin\omega_{d},\cos x_{1},\cdots,\cos x_{d},\sin x_{1},\cdots,\sin x_{d},\mathrm{Re}z,\mathrm{Im}z)>0$ (4.3)
to replace the condition (4.2), where the degree of
$P(\cos\omega_{1},\cdots,\cos\omega_{d},\sin\omega_{1},\cdots,\sin\omega_{d},\cos x_{1},\cdots,\cos x_{d},\sin x_{1},\cdots,\sin x_{d},\mathrm{Re}z,\mathrm{Im}z)$
is at most $C_{1}n_{0}^{2}$. One can further truncate the Taylor series of the trigonometric functions and replace (4.3) by a polynomial inequality
$P(\omega,x,\mathrm{Re}z,\mathrm{Im}z)>0$ (4.4)
whose degree is at most $C_{2}n_{0}^{3}$.
Fix $\omega\in\mathbb{T}^{d}(p,q)$ and $z$. According to Lemma 3.5, the exceptional set $\Omega$ does not only satisfy
$\mathrm{mes}\Omega<\exp(-C_{0}n^{\sigma})$ (4.5)
but also may be assumed semialgebraic of degree less than $Cn_{0}^{3}$. Notice that this set depends on $z$ and $\omega$.
Fix $n=n_{0}$ in Lemma 3.5 and redefine $\Lambda$ to be one of the intervals
$\{[-n_{0}+1,n_{0}-1],[-n_{0},n_{0}-1]\}.$ (4.6)
Let $\Omega=\Omega(z)$ be as above. For $x\notin\Omega$, one of the intervals $\Lambda$ (depending on $x$) satisfies
$|G_{\Lambda}^{\beta,\eta}(m_{1},m_{2};z)|<e^{-\gamma|m_{1}-m_{2}|+n_{0}^{1-}}$ (4.7)
for all $m_{1},m_{2}\in\Lambda$.
Fix $x_{0}\in\mathbb{T}^{d}$. Now we consider the orbit $\{x_{0}+j\omega:|j|\leq n_{1}\}$. In this case, we let $n_{1}=n_{0}^{C}$, where $C$ is a sufficiently large constant. Apply Lemma 3.7 with $\mathcal{S}=\Omega$, $B=C_{2}n_{0}^{3}$, $\eta=\exp(-C_{0}n^{\sigma})$ and $n=n_{1}$. Then the statement (3.15) implies that except for at most $n_{1}^{1-\delta}$ values of $|j|<n_{1}$, taking $x=x_{0}+j\omega$, one of the intervals $\Lambda$ from (4.6) satisfies (4.7).
Assume that $u=\{u(n)\}_{n\in\mathbb{Z}}$ satisfies the equation ~$\tilde{\mathcal{E}}(x_{0},\omega)u=zu$ with $u(0)=1$ and $|u(n)|\lesssim n^{C}$.
If $\Lambda+j=[a,b]$, then $\mathcal{E}_{\Lambda+j}=R_{\Lambda+j}\tilde{\mathcal{E}}R_{\Lambda+j}^{*}$. According to the Poisson’s formula,
$u(n)=$ $G^{\beta,\eta}_{[a,b]}(a,n;z)\begin{cases}(z\overline{\beta}-\alpha_{a})u(a)-\rho_{a}u(a+1),\quad&a\text{ even}\\ (z\alpha_{a}-\beta)u(a)+z\rho_{a}u(a+1),&a\text{ odd}\end{cases}$
$+G^{\beta,\eta}_{[a,b]}(n,b;z)\begin{cases}(z\overline{\eta}-\alpha_{b})u(b)-\rho_{b}u(b-1),\quad&b\text{ even}\\ (z\alpha_{b}-\eta)u(b)+z\rho_{b-1}u(b-1),&b\text{ odd}\end{cases}$
for $a<n<b$. Therefore,
$|u(n)|$ $\leq n^{C}|G^{\beta,\eta}_{[a,b]}(a,n;z)|+n^{C}|G^{\beta,\eta}_{[a,b]}(n,b;z)|$
$\leq n^{C}e^{-\gamma|n-a|+n_{0}^{1-}}+n^{C}e^{-\gamma|b-n|+n_{0}^{1-}}$
$\leq n^{C}e^{n_{0}^{1-}}(e^{-\gamma|n-a|}+e^{-\gamma|b-n|}).$
In particular, taking $n=j$, we have that $|j-a|>\frac{n_{0}}{2}$ and $|j-b|>\frac{n_{0}}{2}$. Then it follows that
$|u(j)|<e^{-\frac{\gamma}{2}n_{0}}$ (4.8)
holds except at most $n^{1-\delta}$ values of $|j|<n_{1}$.
Next, let $I=[-j_{0}+1,j_{0}-1]$. Then $\mathcal{E}_{I}=R_{I}\tilde{\mathcal{E}}R_{I}^{*}$. Hence,
$1=|u(0)|\leq$ $|G^{\beta,\eta}_{I}(-j_{0}+1,0;z)|(2|u(-j_{0}+1)|+|u(-j_{0}+2)|)$
$+|G^{\beta,\eta}_{I}(0,j_{0}-1;z)|(2|u(j_{0}-1)|+|u(j_{0}-2)|).$
If $-j_{0}+1$, $-j_{0}+2$, $j_{0}-2$, $j_{0}-1$ satisfy (4.8), one can obtain that
$\|G_{[-j_{0}+1,j_{0}-1]}^{\beta,\eta}(x_{0},z)\|\geq\frac{1}{6}e^{\frac{\gamma}{2}n_{0}}.$ (4.9)
Equivalently,
$\mathrm{dist}(z,\sigma(\mathcal{E}_{[-j_{0}+1,j_{0}-1]}^{\beta,\eta}(x_{0})))<6e^{-\frac{\gamma}{2}n_{0}}.$ (4.10)
Therefore, fixing $z$, if there is a state $u$ with $u(0)=1$, then for any large $n_{0}$, there exists some $j_{0}$, $|j_{0}|<n_{1}=n_{0}^{C}$ for which (4.10) holds.
Denote $\Sigma_{\omega}=\mathop{\cup}\limits_{|j|\leq n_{1}}\sigma(\mathcal{E}_{-j,j}^{\beta,\eta}(x_{0}))$. It follows from (4.7) and (4.10) that if
$x\notin\mathop{\cup}\limits_{z^{\prime}\in\Sigma_{\omega}}\Omega(z^{\prime}),$ (4.11)
then one of the sets in (4.6) satisfies
$|G_{\Lambda}^{\beta,\eta}(m_{1},m_{2};z)|<e^{-\gamma|m_{1}-m_{2}|+n_{0}^{1-}}$ (4.12)
for $m_{1},m_{2}\in\Lambda$.
Now consider the interval $[\frac{n_{2}}{2},2n_{2}]$ with $n_{2}=n_{0}^{C^{\prime}}$ ($C^{\prime}$ a sufficiently large constant). Suppose that we ensured that
$x_{0}+n\omega\notin\mathop{\cup}\limits_{z^{\prime}\in\Sigma_{\omega}}\Omega(z^{\prime})(\mathrm{mod}\,1)\quad\mathrm{for}\,\,\mathrm{all}\,\,\frac{n_{2}}{2}<|n|<2n_{2}.$ (4.13)
Thus, for each $\frac{n_{2}}{2}<|n|<2n_{2}$, there is an interval
$\Lambda^{(n)}\in\{[-n_{0}+1,n_{0}-1],[-n_{0},n_{0}-1]\}$
for which (4.12) holds:
$|G_{\Lambda^{(n)}+n}^{\beta,\eta}(m_{1},m_{2};z)|<e^{-\gamma|m_{1}-m_{2}|+n_{0}^{1-}}\,\,\mathrm{for}\,\,m_{1},\,m_{2}\in\Lambda^{(n)}+n.$ (4.14)
Define the interval
$\tilde{\Lambda}=\mathop{\cup}\limits_{\frac{n_{2}}{2}<n<2n_{2}}(\Lambda^{(n)}+n)\supset[\frac{n_{2}}{2},2n_{2}].$
Invoking the paving property, one can obtain that
$|G_{\tilde{\Lambda}}^{\beta,\eta}(m_{1},m_{2};z)|<e^{-\frac{\gamma}{2}|m_{1}-m_{2}|+n_{0}^{1-}}\,\,\mathrm{for}\,\,m_{1},\,m_{2}\in[\frac{n_{2}}{2},2n_{2}].$ (4.15)
Therefore, for $k\in[\frac{n_{2}}{2},2n_{2}]$, we get that
$|u(k)|$ $\leq|G_{[\frac{n_{2}}{2},2n_{2}]}^{\beta,\eta}(\frac{n_{2}}{2},k;z)|n_{2}^{C}+|G_{[\frac{n_{2}}{2},2n_{2}]}^{\beta,\eta}(k,2n_{2};z)|n_{2}^{C}$
$<n_{2}^{C}e^{-\frac{\gamma}{2}|k-\frac{n_{2}}{2}|}+n_{2}^{C}e^{-\frac{\gamma}{2}|k-2n_{2}|}$
$<e^{-\frac{\gamma}{2}k}.$
For $k\in\mathbb{Z}_{-}$, we can get the same result. This implies that the exponential decay property follows.
Now we verify the assumption (4.13). For $|j|\leq n_{1}$, we consider the set
=$\mathcal{B}=\{(\omega,z^{\prime},x):\omega\in\mathbb{T}^{d}(p,q),z^{\prime}\in\sigma(\mathcal{E}_{[-j,j]}^{\beta,\eta}(x_{0})),x\in\Omega(z^{\prime})\}\subset\mathbb{T}^{d}\times\partial\mathbb{D}\times\mathbb{T}^{d}.$
For fixed $\omega$, the condition $z^{\prime}\in\sigma(\mathcal{E}_{[-j,j]}^{\beta,\eta}(x_{0}))$ and $x\in\Omega(z^{\prime})$ can be replaced by the inequalities of the polynomials respectively as before. Therefore, $\mathcal{B}$ is a semialgebraic set of degree at most $C_{3}n_{1}^{3}$. Let $\mathcal{S}=\mathrm{Proj}_{(\omega,x)}\mathcal{B}\subset\mathbb{T}^{d}\times\mathbb{T}^{d}$. Obviously, $\mathrm{mes}\mathcal{S}<\exp(-C_{0}n_{1}^{\sigma})$. According to Lemma 3.6, $\mathcal{S}$ is a semialgebraic set whose degree is at most $n_{1}^{3C_{4}}$ for some constant $C_{4}$. Returning to assumption (4.13), we need to verify that
$(\omega,x_{0}+n\omega)\notin\mathcal{S}$ (4.16)
for $\frac{n_{2}}{2}\leq|n|\leq 2n_{2}$. Now we apply Lemma 3.8 with $n=d$, $B=n_{1}^{3C_{4}}$, $\eta=\exp(-C_{0}n_{1}^{\sigma})$. Take $\varepsilon=n_{2}^{-\frac{1}{10}}$. Then there is a decomposition $\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}$. It follows that
$\mathrm{mes}(\mathrm{Proj}_{\omega}\mathcal{S}_{1})<B^{C}\varepsilon<n_{1}^{3C_{4}C}n_{2}^{-\frac{1}{10}}<n_{2}^{-\frac{1}{11}},$ (4.17)
where $n_{2}$ is large enough. Do the partition
$[0,1]^{d}=\mathop{\bigcup}\limits_{l}\Big{(}x_{l}+[0,\frac{1}{n_{2}}]^{d}\Big{)},\,\,l\leq n_{2}^{d}.$
For fixed $\frac{n_{2}}{2}\leq|n|\leq 2n_{2}$ and $l$, we consider $L=\Big{\{}(\omega,x_{0}+nx_{l}+n\omega):\omega\in[0,\frac{1}{n_{2}}]^{d}\Big{\}}$, which is a translate of the $d$-hyperplane $\Big{\{}\frac{1}{n}e_{j}+e_{j+d}:0\leq j\leq d-1\Big{\}}.$ Furthermore, we have that $\mathop{\max}\limits_{0\leq j<d}|\mathrm{Proj}_{L}e_{j}|<\frac{1}{100}\varepsilon<\varepsilon^{2}$. By Lemma 3.8, one can obtain that
$\mathrm{mes}_{\mathbb{T}^{d}}\Big{\{}\omega\in[0,\frac{1}{n_{2}}]^{d}:(\omega,x_{0}+nx_{l}+n\omega)\in\mathcal{S}_{2}\Big{\}}=\mathrm{mes}(\mathcal{S}_{2}\cap L)<B^{C}\varepsilon^{-1}\eta^{\frac{1}{2d}}<n_{1}^{3C_{4}C}n_{2}^{\frac{1}{10}}\exp(-\frac{C_{0}}{2d}n_{1}^{\sigma}).$
Summing the contributions over $n$ and $l$, we have that
$\mathrm{mes}$$\Big{\{}\omega:(\omega,x_{0}+n\omega)\in\mathcal{S}_{2}\,\,\mathrm{for}\,\,\mathrm{some}\,\,\frac{1}{n_{2}}<|n|<2n_{2}\Big{\}}$
$<n_{2}^{d+1}n_{1}^{3C_{4}C}n_{2}^{\frac{1}{10}}\exp(-\frac{C_{0}}{2d}n_{1}^{\sigma})<\exp(-n_{1}^{\frac{\sigma}{2}}).$ (4.18)
From (4.17) and (4.18), we can get an $\omega$-set whose measure is at most
$n_{2}^{-\frac{1}{11}}+\exp(-n_{1}^{\frac{\sigma}{2}})<n_{2}^{-\frac{1}{12}}.$
Summing these sets over $j$, $j\leq n_{1}$, we get an $\omega$-set of measure at most $n_{1}n_{2}^{-\frac{1}{12}}<n_{2}^{-\frac{1}{13}}$. This implies that (4.13) holds for $\omega$ outside the $\omega$-set we obtained. Since we initially fixed $n_{0}$ and set $n_{1}=n_{0}^{C}$, $n_{2}=n_{0}^{C^{\prime}}$, we can denote the above $\omega$-set to be $\Omega_{n_{0}}$. Then we have that
$\mathrm{mes}\Omega_{n_{0}}<n_{2}^{-\frac{1}{13}}=n_{0}^{-\frac{C^{\prime}}{13}}<n_{0}^{-10}.$
According to the above analysis, we ensure that if
$\mathcal{E}(x_{0},z)u=zu,\,\,u(0)=1,\,\,|u(n)|\lesssim n^{C},$
then
$|u(k)|<e^{-\frac{\gamma}{4}k}\quad\mathrm{for}\,\,\,k\in\big{[}\frac{n_{2}}{2},2n_{2}\big{]}.$
Let $\Omega_{*}=\mathop{\cup}\limits_{n}\mathop{\cap}\limits_{n_{0}>n}\Omega_{n_{0}}$. It follows that $\mathrm{mes}\Omega_{*}=0$ and Anderson localization holds for $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{*}$.
[…]
5 Quantum walks
Quantum walks share a structural similarity with CMV matrices via unitary equivalence. Now we recall some basic knowledge about quantum walks, see [9, Section 2.4]. A quantum walk is described by a unitary operator on the Hilbert space $\mathcal{H}=\ell^{2}(\mathbb{Z})\otimes\mathbb{C}^{2}$, which models a state space in which a wave packet comes equipped with a “spin” at each integer site. Notice that the elementary tensors of the form $\delta_{n}\otimes e_{\uparrow}$ and $\delta_{n}\otimes e_{\downarrow}$ with $n\in\mathbb{Z}$ comprise an orthonormal basis of $\mathcal{H}$. Here, $\{e_{\uparrow},e_{\downarrow}\}$ denotes the canonical basis of $\mathbb{C}^{2}$. A time-homogeneous quantum walk scenario is given as soon as coins
$C_{n}=\left[\begin{array}[]{cc}c_{n}^{11}&c_{n}^{12}\\ c_{n}^{21}&c_{n}^{22}\end{array}\right]\in\mathbb{U}(2),\quad n\in\mathbb{Z}$ (5.1)
are specified. Assume that $c_{n}^{11},c_{n}^{22}\neq 0$. As one passes from time $t$ to $t+1$, the update rule of the quantum walk is
$\delta_{n}\otimes e_{\uparrow}\mapsto c_{n}^{11}\delta_{n+1}\otimes e_{\uparrow}+c_{n}^{21}\delta_{n-1}\otimes e_{\downarrow},$
$\delta_{n}\otimes e_{\downarrow}\mapsto c_{n}^{12}\delta_{n+1}\otimes e_{\uparrow}+c_{n}^{22}\delta_{n-1}\otimes e_{\downarrow}.$
If we extend this by linearity and continuity to general elements of $\mathcal{H}$, this defines a unitary operator $U$ on $\mathcal{H}$. For a typical element $\psi\in\mathcal{H}$, we may describe the action of $U$ in coordinates via
$[U\psi]_{\uparrow,n}=c_{n-1}^{11}\psi_{\uparrow,n-1}+c_{n-1}^{12}\psi_{\downarrow,n-1},$
$[U\psi]_{\downarrow,n}=c_{n+1}^{21}\psi_{\uparrow,n+1}+c_{n+1}^{22}\psi_{\downarrow,n+1}.$
Then the matrix representation of $U$ is
$U=\left[\begin{array}[]{ccccccccc}\ddots&\ddots&\ddots&\ddots&&&&&\\ &0&0&c_{0}^{21}&c_{0}^{22}&&&&\\ &c_{-1}^{11}&c_{-1}^{12}&0&0&&&&\\ &&&0&0&c_{1}^{21}&c_{1}^{22}&&\\ &&&c_{0}^{11}&c_{0}^{12}&0&0&&\\ &&&&&0&0&c_{2}^{21}&c_{2}^{22}\\ &&&&&c_{1}^{11}&c_{1}^{12}&0&0\\ &&&&&\ddots&\ddots&\ddots&\ddots\end{array}\right].$ (5.2)
Comparing the above matrix with the extended CMV matrix, the structures of these two matrices are similar. If all Verblunsky coefficients with even index vanish, then the extended CMV matrix becomes
$\mathcal{E}=\left[\begin{array}[]{ccccccccc}\ddots&\ddots&\ddots&\ddots&&&&&\\ &0&0&\bar{\alpha}_{1}&\rho_{1}&&&&\\ &\rho_{-1}&-\alpha_{-1}&0&0&&&&\\ &&&0&0&\bar{\alpha}_{3}&\rho_{3}&&\\ &&&\rho_{1}&-\alpha_{1}&0&0&&\\ &&&&&0&0&\bar{\alpha}_{5}&\rho_{5}\\ &&&&&\rho_{3}&-\alpha_{3}&0&0\\ &&&&&\ddots&\ddots&\ddots&\ddots\end{array}\right].$ (5.3)
The matrix in (5.3) strongly resembles the matrix representation of $U$ in (5.2). However, $\rho_{n}>0$ for all $n$, then (5.2) and (5.3) may not match exactly when $c_{n}^{kk}$ is not real and positive. Indeed, this can be easily resolved by conjugation with a suitable diagonal unitary, as shown in [CMV-quantum-walks].
Given $U$ as in (5.1), write
$c_{n}^{kk}=r_{n}\omega_{n}^{k},\,\,\,n\in\mathbb{Z},\,\,\,k\in\{1,2\},\,\,\,r_{n}>0,\,\,\,|\omega_{n}^{k}|=1.$
Define $\{\lambda_{n}\}_{n\in\mathbb{Z}}$ as
$\lambda_{0}=1,\,\,\,\lambda_{-1}=1,\,\,\,\lambda_{2n+2}=\omega_{n}^{1}\lambda_{2n},\,\,\,\lambda_{2n+1}=\overline{\omega_{n}^{2}}\lambda_{2n-1}.$
Let $\Lambda=\mathrm{diag}(\cdots,\lambda_{-1},\lambda_{0},\lambda_{1},\cdots)$. Then one can obtain that
^$\hat{\mathcal{E}}=\Lambda^{*}U\Lambda,$
where $\hat{\mathcal{E}}$ is the extended CMV matrix whose Verblunsky coefficients are defined by
$\hat{\alpha}_{2n}=0,\,\,\,\hat{\alpha}_{2n+1}=\frac{\lambda_{2n-1}}{\lambda_{2n}}\cdot\overline{c_{n}^{21}}=-\frac{\lambda_{2n+1}}{\lambda_{2n+2}}\cdot c_{n}^{12},\,\,\,n\in\mathbb{Z}.$ (5.4)
Notice that the hypotheses $c_{n}^{11},c_{n}^{22}\neq 0$ imply $\rho_{n}>0$ for all $n\in\mathbb{Z}$. According to the analysis above, there is a close relation between a quantum walk and an extended CMV matrix whose Verblunsky coefficients with even index are zero.
5.1 Gesztesy-Zinchenko cocycle
Now we consider the case of coined quantum walks on the integer lattice where the coins are distributed quasi-periodically, i.e.,
$C_{n}=C_{n,x,\omega}=\left[\begin{array}[]{cc}c_{n}^{11}(x+n\omega)&c_{n}^{12}(x+n\omega)\\ c_{n}^{21}(x+n\omega)&c_{n}^{22}(x+n\omega)\end{array}\right]$
with analytic sampling functions and $\omega\notin\mathbb{Q}$. We denote the corresponding quantum walk by $U_{\omega}(x)$.
As for the extended CMV matrix $\hat{\mathcal{E}}$, there is a decomposition $\hat{\mathcal{E}}=\hat{\mathcal{L}}\hat{\mathcal{M}}$ (see section 3.4 for details). For a given solution $u$ of ^$\hat{\mathcal{E}}u=zu$, we define $v=\hat{\mathcal{L}}^{-1}u$. One can verify that ^$\hat{\mathcal{M}}u=zv$. Then according to [13, section 2], we have that
$\left[\begin{array}[]{c}u(n+1)\\ v(n+1)\end{array}\right]=Y(n,z)\left[\begin{array}[]{c}u(n)\\ v(n)\end{array}\right],$
where
$Y(n,z)=\frac{1}{\rho_{n}}\begin{cases}\left[\begin{array}[]{cc}-\hat{\alpha}_{n}&1\\ 1&-\overline{\hat{\alpha}}_{n}\end{array}\right]\quad n\,\,\text{is even},\\ \\ \left[\begin{array}[]{cc}-\overline{\hat{\alpha}}_{n}&z\\ z^{-1}&-\hat{\alpha}_{n}\end{array}\right]\quad n\,\,\text{is odd}.\end{cases}\$
For the sake of convenience, we write $Y(n,z)=P(\hat{\alpha}_{n},z)$ when $n$ is even and $Y(n,z)=Q(\hat{\alpha}_{n},z)$ when $n$ is odd. According to [8, section 3], we can define Gesztesy-Zinchenko cocycle by
$G(\omega,z;x)=Q(\hat{\alpha}(x+\omega),z)P(\hat{\alpha}(x),z).$
Based upon the above discussion, if the Verblunsky coefficients with even index are equal to zero, then the Gesztesy-Zinchenko cocycle is
$G(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}z&-\overline{\hat{\alpha}}(x)\\ -\hat{\alpha}(x)&z^{-1}\end{array}\right].$
Correspondingly, the $n$-step transfer matrix is
$G_{n}(\omega,z;x)=\prod_{j=n-1}^{0}G(\omega,z;x+j\omega).$
Therefore, the associated Lyapunov exponent of the quantum walk $U_{\omega}(x)$ can be defined by
$L^{U}(w,z):=\lim_{n\rightarrow\infty}\frac{1}{n}\int_{\mathbb{T}^{d}}\log\|G_{n}(\omega,z;x)\|dx.$
5.2 Anderson localization for the quantum walks
According to the analysis in section 2, the Szegő cocycle and the $n$-step transfer matrix of the matrix $\hat{\mathcal{E}}$ are
$\hat{M}(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\hat{\alpha}}(x)}{\sqrt{z}}\\ -\hat{\alpha}(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right]$
and $\hat{M}_{n}(\omega,z;x)=\mathop{\prod}\limits_{j=n-1}^{0}\hat{M}(\omega,z;x+j\omega)$, respectively. In addition, the Lyapunov exponent of the matrix $\hat{\mathcal{E}}$ is
$\hat{L}(w,z):=\lim_{n\rightarrow\infty}\frac{1}{n}\int_{\mathbb{T}^{d}}\log\|\hat{M}_{n}(\omega,z;x)\|dx.$
By a simple calculation, we know that
$G(\omega,z;x)=\left[\begin{array}[]{cc}\frac{1}{\sqrt{z}}&0\\ 0&\sqrt{z}\end{array}\right]\hat{M}(\omega,z;x).$
It is obvious that the difference between the Lyapunov exponent of the quantum walk $U_{\omega}(x)$ and the extended CMV matrix $\hat{\mathcal{E}}$ depends on the norm of the matrix
$\left[\begin{array}[]{cc}\frac{1}{\sqrt{z}}&0\\ 0&\sqrt{z}\end{array}\right]=:M_{0}.$
Let $C_{\mathrm{nor}}:=\|M_{0}\|$. Then one can obtain that
$L^{U}(w,z)$ $=\lim_{n\rightarrow\infty}\frac{1}{n}\int_{\mathbb{T}^{d}}\log\|G_{n}(\omega,z;x)\|dx$
[…]
$=\lim_{n\rightarrow\infty}\frac{1}{n}\int_{\mathbb{T}^{d}}\log(C_{\mathrm{nor}})^{n}+\log\|\hat{M}_{n}(\omega,z;x)\|)dx$
$=\log C_{\mathrm{nor}}+\hat{L}(w,z).$
On the other hand, when $z=1$, $G_{n}(\omega,z;x)=\hat{M}_{n}(\omega,z;x)$. It follows that
$L^{U}(w,z)\geq\hat{L}(w,z).$
From the above discussion, quantum walks can be viewed as a special class of extended CMV matrices whose Veblunsky coefficients with even index are equal to zero. Due to the relationship between the quantum walks and the extended CMV matrices, we can get the Anderson localization for the quantum walks with multi-frequency quasi-periodic coins.
Theorem 5.1.
The family of multi-frequency quasi-periodic quantum walks $\{U_{\omega}(x)\}$ with analytic coins display Anderson localization in the regime of positive Lyapunov exponents for a fixed phase vector $x_{0}$ and almost all frequency vectors satisfying the standard Diophantine condition.
Remark 5.2.
This statement is an application of Theorem 1.1. Since
$\hat{L}(w,z)\leq L^{U}(w,z)\leq\log C_{\mathrm{nor}}+\hat{L}(w,z),$
then the Anderson localization result for the quantum walks follows.
$\mathbf{Acknowledgments}$
This work was supported in part by the NSFC (No. 11571327, 11971059).